# Discovering Effective Policies for Land-Use Planning with Neuroevolution

### v5

@@ Abstract.
How areas of land are allocated for different uses, such as forests, urban areas, and agriculture, has a large effect on the terrestrial carbon balance, and therefore climate change. Based on available historical data on land-use changes and a simulation of the associated carbon emissions and removals, a surrogate model can be learned that makes it possible to evaluate the different options available to decision-makers efficiently. An evolutionary search process can then be used to discover effective land-use policies for specific locations. Such a system was built on the Project Resilience platform (ITU, 2023) and evaluated with the Land-Use Harmonization dataset LUH2 (Hurtt et al., 2020) and the bookkeeping model BLUE (Hansis et al., 2015). It generates Pareto fronts that trade off carbon impact and amount of land-use change customized to different locations, thus providing a potentially useful tool for land-use planning.
Key words and phrases:
Surrogate modeling, Decision making, Climate Change, Land-use Change, Neural Networks, Evolutionary Computation<ccs2012> <concept> <concept_id>10010147.10010257.10010321</concept_id> <concept_desc>Computing methodologies Machine learning algorithms</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010919.10010172</concept_id> <concept_desc>Computing methodologies Distributed algorithms</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10011809.10011812</concept_id> <concept_desc>Computing methodologies Genetic algorithms</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10010294</concept_id> <concept_desc>Computing methodologies Neural networks</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
1. Introduction
One of the main factors contributing to climate change is CO2 emissions due to land-use change (Friedlingstein et al., 2023). Land-use CO2 emissions crucially depend on how much land area is allocated for different uses. Forests in general take up more carbon than e.g. croplands and rangelands, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon removals while maintaining economic viability.
@@ 1. Introduction
An approach for land-use optimization was developed as part of Project Resilience, a non-profit project hosted by the ITU agency of the United Nations (ITU, 2023). The goal is to provide decision-makers with a tool to know how their land-use choices affect CO2 fluxes in the long term, and make suggestions for optimizing these choices.
@@
More specifically, the tool is designed to answer three questions: (1) For a geographical grid cell, identified by its latitude and longitude, what changes to the land use can be made to reduce CO2 emissions? (2) What will be the long-term CO2 impact of changing land use in a particular way? (3) What are the optimal land-use choices that can be made with minimal cost and maximal effect?
@@
The approach is based on the Evolutionary Surrogate-assisted Prescription method (Francon et al., 2020). The idea is to first utilize historical data to learn a surrogate model on how land-use decisions in different contexts affect carbon emissions, and then use this model to evaluate candidates in an evolutionary search process for good land-use change policies.
@@
As a result, a Pareto front is generated of solutions that trade off reduction in carbon emissions and the amount of change in land use. Each point in the Pareto front represents an optimal policy for that tradeoff. To make the results trustworthy, the tool allows the decision-maker to explore modifications to these policies, and see the expected effect. In the future, it should also be possible to evaluate confidence of the predictions, evolve rule sets to make the policies explainable, and utilize ensembling and further objectives and preferences to make them more accurate. Thus, the tool harnesses several techniques in machine learning to provide a practical tool for decision-makers in optimizing land-use decisions.
@@
Project code and data are available at https://github.com/Project-Resilience/mvp/tree/main/use_cases/eluc, and an interactive demo at https://landuse.evolution.ml.
2. Background
@@
2.1. Evolutionary Surrogate-assisted Prescription
Evolutionary Surrogate-assisted Prescription (ESP; Francon et al., 2020) is an approach for optimizing decision-making in a variety of domains (Figure 1). The main idea is that a decision policy can be represented as a neural network, or a set of rules, and a good policy can be discovered through population-based search, i.e. using evolutionary computation techniques. However, each candidate must be evaluated, which is difficult to do in many real-world applications. Therefore, a surrogate model of the world is learned from historical data, predicting how good the resulting outcomes are for each decision in each context.
@@ 2.1. Evolutionary Surrogate-assisted Prescription
such that over all possible contexts and outcome dimensions is maximized (assuming increased outcomes are better). It thus approximates the optimal decision policy for the problem. The predictor can be learned from historical data on triples. In contrast, the optimal actions for each context are not known, and must therefore be found through search.
@@
ESP was first evaluated in reinforcement learning tasks such as function approximation, cart-pole control, and the flappybird game, and found to discover significantly better solutions, find them faster, and with lower regret than standard approaches such as direct evolution, Deep Q-Networks (DQN; Mnih et al., 2015), and Proximal Policy Optimization (PPO; Schulman et al., 2017). Its power comes from automatic regularization and automatic curricular learning: When predictors are learned at the same time as the prescriptors, each prescriptor is evaluated against multiple predictors; as the predictors improve, they provide more refined evaluations. ESP is thus a powerful mechanism for learning policies in time-varying domains.
2.2. Project Resilience
@@
A major application of ESP to decision-making was developed for optimizing strategies for non-pharmaceutical interventions (NPIs) in the COVID-19 pandemic (Miikkulainen et al., 2021). Using case and NPI data from around the world, collected in 2020-2022 by Oxford University’s COVID-19 project (Hale et al., 2020), a neural network model was developed to predict the course of the pandemic; this model was then used as a surrogate in evolving another neural network to prescribe NPIs. It discovered different strategies at different phases of the pandemic, such as focusing on schools and workplaces early on, alternating policies over time, and focusing on public information and masking later. This application demonstrated the power of ESP in discovering good tradeoffs for complex real-world decision-making tasks.
@@
Encouraged by this application, the XPRIZE Pandemic Response Challenge (XPRIZE, 2023) was developed in 2020-2021 for NPI optimization using the same predictor+prescriptor approach. The competitors used several different machine learning or other techniques, and drew significant attention to AI-supported decision-making. The results were effective and in some cases used to inform decision-makers (e.g. on the 2021 surge in Valencia, Spain, and on the fall 2021 school openings in Iceland). An important conclusion was that it was possible to convene a group of experts—data scientists, epidemiologists, and public health officials—to build a useful set of tools to advise decision-makers on how to cope with and plan around a health disruption to society.
@@
This experience with the XPRIZE Pandemic Response Challenge then led to Project Resilience (ITU, 2023). The XPRIZE competition serves as a blueprint for a collaborative effort to tackle global problems. The goal of Project Resilience is to build a public AI utility where a global community of innovators and thought leaders can enhance and utilize a collection of data and AI approaches to help with better preparedness, intervention, and response to environmental, health, information, social equity, and similarly scoped problems. The project is hosted by the ITU agency of the United Nations, under its Global Initiative for AI and Data Commons, and contributes to its general efforts toward meeting the Sustainable Development Goals (SDGs).
@@
3. Land-Use Optimization Task
@@
3.1. Data
The data for carbon emissions (Emissions resulting from Land-Use Change, ELUC) originate from the Bookkeeping of Land-Use Emissions model (BLUE; Hansis et al., 2015), which attributes carbon fluxes to land-use activities. BLUE is one of three bookkeeping models that are used to provide ELUC estimates for the annually published Global Carbon Budget (GCB; Friedlingstein et al., 2023). Out of these three models, BLUE is the only one that provides spatially explicit ELUC estimates. In this paper, BLUE is used to estimate committed emissions due to land-use changes, which means that all the emissions or removals caused by a land-use change event are attributed to the year of the event. The emission (or removal) estimates thus relate to the long-term impact of a land-use change event, i.e. up to multiple decades (Hansis et al., 2015; Houghton et al., 1983). While in principle BLUE can be used as the surrogate model for ESP, in practice its calculations are too expensive to carry out on demand during the search for good policies. Therefore, we performed a number of BLUE simulations covering a comprehensive set of land-use changes for 1850-2022, resulting in a dataset that could be used to train an efficient surrogate model.
The Land-Use Change (LUC) data is provided by the Land-Use Harmonization project (LUH2; Hurtt et al., 2020). LUH2 provides data on fractional land-use patterns, underlying land-use transitions, and key agricultural management information, annually for the time period 850-2100 at 0.250.25 degree resolution. LUH2 provides separate layers for states, transitions, and management. Version LUH2-GCB2022 from the Global Carbon Budget 2022 (Friedlingstein et al., 2022) was the basis for this paper. It is based on the HYDE3.3 historical database Klein Goldewijk et al. (2017); Kees Klein Goldewijk and van Zanden (2017) and uses the 2021 wood harvest data provided by the FAO agency of the United Nations. It consists of the following land-use types:
Primary: Vegetation that is untouched by humans
@@ 3.1. Data
Secondary: Vegetation that has been touched by humans
@@
crop: All types pooled into one category
@@
BLUE calculates yearly ELUC estimates based on the land-use changes provided by LUH2. Urban areas are assumed to have the same carbon content as croplands. All crop types are currently pooled into a single category. BLUE considers that the degradation of primary to secondary land often lowers the carbon content in vegetation and soils, and it considers pastures and the degradation of natural grasslands and savannahs by grazing. A complete overview of the model can be found in Hansis et al. (2015).
@@
3.2. Decision-making Problem
The modeling approach aims to understand the domain in two ways: (1) In a particular situation, what are the outcomes of the decision maker’s actions? (2) What are the actions that result in the best outcomes, i.e. the lowest carbon emission (or highest removals) with the smallest change in land use? The data is thus organized into context, action, and outcome variables.
@@ 3.2. Decision-making Problem
Latitude and Longitude, representing the cell on the grid.
@@
Land use, representing the percentage of land in the cell used by each type above. For simplicity, all crop types were pooled into a single type. In addition, a ‘nonland’ type represents the percentage of the cell that is not land (e.g. typically sea, lake, etc.). The percentages sum up to 100% in each grid cell.
@@
First, decision-makers cannot affect primary land. The idea is that it is always better to preserve primary vegetation, such as pristine rainforest; destroying it is not an option given to the system. Technically, it is not possible to re-plant primary vegetation. Once destroyed, it is destroyed forever. If re-planted, it would become secondary vegetation, for which BLUE assumes lower carbon content than for primary vegetation Hansis et al. (2015).
Second, decision-makers cannot affect urban areas. The needs of urban areas are dictated by other imperatives, and optimized by other decision makers. Therefore, the ESP system cannot recommend that a city should be destroyed, or expanded.
Outcomes consist of two conflicting variables. The primary variable is ELUC, i.e. emissions from land-use change. It consists of all CO2 emissions attributed to the land-use change, in metric tons of carbon per hectare (tC/ha), obtained from the BLUE simulation. Note that all follow-up CO2 emissions (or removals) are attributed to the year of the land-use change, thus neglecting temporal dynamics in CO2 fluxes. A positive number means carbon is emitted to the atmosphere and a negative number means carbon is taken up by vegetation. The secondary variable is the cost of the land-use change, represented by the percentage of land that was changed. This variable is calculated directly from the actions.

### v7

@@ Abstract
How areas of land are allocated for different uses, such as forests, urban areas, and agriculture, has a large effect on the terrestrial carbon balance, and therefore climate change. Based on available historical data on land-use changes and a simulation of the associated carbon emissions and removals, a surrogate model can be learned that makes it possible to evaluate the different options available to decision-makers efficiently. An evolutionary search process can then be used to discover effective land-use policies for specific locations. Such a system was built on the Project Resilience platform [17] and evaluated with the Land-Use Harmonization dataset LUH2 [16] and the bookkeeping model BLUE [14]. It generates Pareto fronts that trade off carbon impact and amount of land-use change customized to different locations, thus providing a proof-of-concept tool that is potentially useful for land-use planning.
1 Introduction
One of the main factors contributing to climate change is CO2 emissions due to land-use change [8]. Land-use CO2 emissions depend crucially on how much land area is allocated for different uses. Certain land-use changes such as afforestation can absorb carbon, and are considered a negative emission technology (NET) [24]. However not all land can be simply converted to forest; for example, land used for crops and pasture is essential for other purposes such as food supply. Land-use patterns must therefore be planned to minimize carbon emissions while maintaining a satisfactory level of benefits received from the land, or ecosystem services. It is also important to note that land-use planning on its own is not enough to reach net-zero emissions. However, it can help offset emissions from sectors that are harder to decarbonize.
Recently, several different approaches were taken in order to optimize land-use planning to minimize carbon emissions. One such approach did so for the city of Nanjing in China using multi-objective linear programming [3]. While effective, this optimization method cannot take advantage of nonlinearities. In another approach, multi-criteria decision-making (MCDM) was used to optimize urban development in the Rajshahi Metropolitan Area of Bangladesh [29]. This approach uses manually ordered weights on outcomes to create solutions that trade off multiple objectives [37]. Six pre-defined scenarios with fixed levels of trade-off are required in order to construct such weights. The land-use planning task was also represented as a causal machine learning task in the potential outcomes framework [31] with crop type as the treatment, environmental factors as the features, and net primary productivity (NPP, i.e. the uptake flux in which carbon from the atmosphere is sequestrated by plants) as the outcome [12]. Causality between treatment and outcome was found to inform policy, but the approach does not create a comprehensive strategy. Thus, all three methods were effective in their approach but each had their own unique drawbacks.
@@ 1 Introduction
In this paper, a new approach for land-use optimization was developed with these concerns in mind. It takes advantage of nonlinearities in the data and automatically generates a Pareto front of land-use planning policies that trade off multiple outcomes. This work was done as part of Project Resilience, a non-profit project hosted by the ITU agency of the United Nations [17]. The goal is to provide decision-makers with a tool to know how their land-use choices affect CO2 fluxes in the long term, and make suggestions for optimizing these choices.
@@
More specifically, the tool was designed to answer three questions: (1) For a geographical grid cell identified by its latitude and longitude, what changes to the land use can be made to reduce overall CO2 emissions (or in some cases, maximize carbon absorption)? (2) What will be the long-term CO2 impact of changing land use in a particular way? (3) What are the optimal land-use choices that can be made with minimal cost and maximal effect?
@@
The approach is based on the Evolutionary Surrogate-assisted Prescription method [7]. The idea is to first utilize historical data to learn a surrogate model on how land-use decisions in different contexts affect carbon emissions from land-use change, and then use this model to evaluate candidates in an evolutionary search process for good land-use change policies.
@@
As a result, a Pareto front is generated of solutions that trade off reduction in carbon emissions and the amount of change in land use. Each point in the Pareto front represents an optimal policy for that trade-off. To make the results trustworthy, the tool allows the decision-maker to explore modifications to these policies, and see the expected effect. In the future, it should also be possible to evaluate confidence of the predictions, evolve rule sets to make the policies explainable, and utilize ensembling and further objectives and preferences to make them more accurate. Thus, the tool harnesses several techniques in machine learning to provide a proof-of-concept for decision-makers in optimizing land-use decisions.
@@
The project page can be found at https://project-resilience.github.io/platform/projects/landuse.html, and an interactive demo at https://landuse.evolution.ml.
2 Background
@@
2.1 Evolutionary Surrogate-assisted Prescription
Evolutionary Surrogate-assisted Prescription [ESP; 7] is an approach for optimizing decision-making in a variety of domains (Figure 1). The main idea is that a decision policy can be represented as a neural network (or in some cases, as a set of rules), and a good policy can be discovered through population-based search, i.e. using evolutionary computation techniques, in order to prescribe actions in a given context. However, each candidate must be evaluated, which is difficult to do in many real-world applications. Therefore, a surrogate model of the world is learned from historical data, predicting how good the resulting outcomes are for each decision in each context.
@@ 2.1 Evolutionary Surrogate-assisted Prescription
such that over all possible contexts and outcome dimensions is minimized (assuming decreased outcomes are better). It thus approximates the optimal decision policy for the problem. The predictor can be learned from historical data on triples. In contrast, the optimal actions for each context are not known, and must therefore be found through search.
More specifically in land-use optimization, context refers to the details of a given area such as current land use, a decision policy generates changes to the land use for the given area based on these details, and the outcomes are (1) the emissions from land-use change (ELUC) predicted by the surrogate as a result of this change, and (2) land-cost change, which is calculated as the percentage of the area changed.
@@
ESP was first evaluated in reinforcement learning tasks such as function approximation, cart-pole control, and the Flappy Bird game, and found to discover significantly better solutions, find them faster, and with lower regret (deviation from the optimal reward) than standard approaches such as direct evolution, Deep Q-Networks [DQN; 23], and Proximal Policy Optimization [PPO; 33]. It thus forms a promising foundation for optimizing land use as well.
2.2 Project Resilience
@@
A major application of ESP to decision-making was developed for optimizing strategies for non-pharmaceutical interventions (NPIs) in the COVID-19 pandemic [22]. Using case and NPI data from around the world, collected in 2020-2022 by Oxford University’s COVID-19 project [13], a neural network model was developed to predict the course of the pandemic; this model was then used as a surrogate in evolving another neural network to prescribe NPIs. It discovered different strategies at different phases of the pandemic, such as focusing on schools and workplaces early on, alternating policies over time, and focusing on public information and masking later. This application demonstrated the power of ESP in discovering good tradeoffs for complex real-world decision-making tasks.
@@
Encouraged by this application, the XPRIZE Pandemic Response Challenge [36] was developed in 2020-2021 for NPI optimization using the same predictor+prescriptor approach. The competitors used several different machine learning or other techniques, and drew significant attention to AI-supported decision-making. The results were effective and in some cases used to inform decision-makers (e.g. on the 2021 surge in Valencia, Spain, and on the fall 2021 school openings in Iceland). An important conclusion was that it was possible to convene a group of experts—data scientists, epidemiologists, and public health officials—to build a useful set of tools to advise decision-makers on how to cope with and plan around a health disruption to society.
@@
This experience with the XPRIZE Pandemic Response Challenge then led to Project Resilience [17]. The XPRIZE competition serves as a blueprint for a collaborative effort to tackle global problems. The goal of Project Resilience is to build a public AI utility where a global community of innovators and thought leaders can enhance and utilize a collection of data and AI approaches to help with better preparedness, intervention, and response to environmental, health, information, social equity, and similarly scoped problems. The project is hosted by the ITU agency of the United Nations, under its Global Initiative for AI and Data Commons, and contributes to its general efforts toward meeting the Sustainable Development Goals (SDGs).
@@
3 Land-Use Optimization Task
@@
3.1 Data
The data for carbon emissions (Emissions resulting from Land-Use Change, ELUC) originate from the Bookkeeping of Land-Use Emissions model [BLUE; 14], which attributes carbon fluxes to land-use activities. A bookkeeping model is a semi-empirical statistical model that calculates CO2 fluxes based on changes in the carbon content of biomass and soil caused by land-use changes. Bookkeeping models employ different carbon densities for different vegetation types (plant functional types in BLUE) and use response functions to determine the temporal dynamics of the fluxes. Advantages of bookkeeping models include their ability to separate anthropogenic from natural CO2 fluxes and to attribute fluxes to specific land-use change events [6]. Bookkeeping models thus differ from more general climate models and land surface models that are based on parameterizations of physical, chemical, and biological processes. BLUE is one of four bookkeeping models that are used to provide ELUC estimates for the annually published Global Carbon Budget [GCB; 9, 8, 10]. BLUE is particularly useful because it provides spatially explicit ELUC estimates. In this paper, BLUE is used to estimate committed emissions due to land-use changes, which means that all the emissions or removals caused by a land-use change event are attributed to the year of the event. The emission (or removal) estimates thus relate to the long-term impact of a land-use change event, i.e. up to multiple decades [14, 15]. While in principle BLUE can be used as the surrogate model for ESP, in practice its calculations are too expensive to carry out on demand during the search for good policies. Therefore, a number of BLUE simulations were performed covering a comprehensive set of land-use changes for 1850-2022, resulting in a dataset that could be used to train an efficient surrogate model.
The Land-Use Change (LUC) data is provided by the Land-Use Harmonization project [LUH2; 16]. LUH2 provides data on fractional land-use patterns, underlying land-use transitions, and key agricultural management information, annually for the time period 850-2100 at 0.250.25 degree resolution. LUH2 provides separate layers for states, transitions, and management. Version LUH2-GCB2022 from the Global Carbon Budget 2022 [9] was the basis for this paper. It is based on the HYDE3.3 historical database [19, 18] and uses the 2021 wood harvest data provided by the FAO agency of the United Nations. It consists of the following land-use types:
Primary: Vegetation that has never been impacted by human activities (e.g. agriculture or wood harvesting) since the beginning of the simulation
@@ 3.1 Data
Secondary: Vegetation that is recovering from previous human disturbance (either wood harvesting or agricultural abandonment)
@@
crop: All crop types pooled into one category
@@
BLUE calculates yearly ELUC estimates based on the land-use changes provided by LUH2. Urban areas are pooled into a single type; their carbon contents are too low and the total urban area on the planet is too small for differences among them to have a significant effect on the CO2 flux. All crop types are also pooled into a single crop category. BLUE considers that the degradation of primary to secondary land often lowers the carbon content in vegetation and soils, and it considers pastures and the degradation of natural grasslands and savannahs by grazing. A complete overview of the model can be found in [14].
It is important to note two caveats in this data. First, the 0.25-degree spatial resolution of LUH2 is rather coarse for land-use recommendations. Second, different crop types do have noticeable differences when making decisions at a local level. To address these concerns, in the future more fine-grained land-use estimates can be used and crop type can be added as a feature variable, and such extensions would allow for more accurate prescriptions at the local level. In the meantime, the current versions of LUH2 and BLUE are sufficient to build a proof-of-concept system that demonstrates the power of land-use planning through neuroevolution.
@@
3.2 Decision-making Problem
The modeling approach aims to understand the domain in two ways: (1) In a particular situation, what are the outcomes of the decision maker’s actions? (2) What are the actions that result in the best outcomes, i.e. the lowest ELUC with the smallest change in land use? The data is thus organized into context, action, and outcome variables.
@@ 3.2 Decision-making Problem
Latitude and Longitude, representing the cell on the grid. Climate information loosely corresponding to plant functional types is implicitly represented in the location of the cell, as shown in Appendix A.
@@
Land use, representing the percentage of land in the cell used by each type above. For simplicity, all crop types were pooled into a single type. The percentages do not necessarily reach 100% in each grid cell; a ’nonland’ type (e.g. typically sea, lake, etc.) is assumed to fill the remainder.
@@
First, decision-makers cannot affect primary land. The idea is that it is always better to preserve primary vegetation, such as pristine rainforests; destroying it is not an option given to the system. Technically, it is not possible to re-plant primary vegetation. Once destroyed, it is destroyed forever. If re-planted, it would become secondary vegetation, for which BLUE assumes lower carbon content than for primary vegetation [14].
Second, decision-makers cannot affect urban areas. The needs of urban areas are dictated by other imperatives, and optimized by other decision makers. Therefore, the ESP system cannot recommend that a city should be destroyed, or expanded. Further constraints on crop usage for food production are discussed in Section 5.4
Outcomes consist of two conflicting variables. The primary variable is ELUC, i.e. emissions from land-use change. It consists of all CO2 emissions attributed to the land-use change, in metric tons of carbon per hectare (tC/ha), obtained from the BLUE simulation. A positive ELUC means carbon is emitted to the atmosphere and a negative number means carbon is taken up by vegetation. For simplicity, all future CO2 emissions or removals are attributed to the initial year of the land-use change; temporal dynamics in CO2 fluxes are not taken into account. Similarly, the full land-use change is assumed to take place; incremental changes are not taken into account. The secondary variable is the cost of the land-use change, represented by the percentage of land that was changed. This variable is calculated directly from the actions.
@@
4 Models
The system consists of the predictor, trained with supervised learning on the historical data, and the prescriptor, discovered through a evolutionary search in which each candidate is evaluated with the predictor.
Prediction: Given the context and actions that were performed, the predictive model estimates the outcomes. In this case, since the cost outcome (i.e. the percentage of changed land) can be calculated directly, only the ELUC is predicted by the model. That is, given the land use of a specific location, and the land-use changes that were made during a specific year, the model predicts the long-term CO2 emissions directly caused by these land-use changes. The model does not predict non-ELUC emissions such as forest fires that may lead to a change into secondary forest.
Any predictive model can be used in this task, including linear regression, random forest, or a neural network; they are each evaluated below. As usual in machine learning, each model is fit to the existing historical data and evaluated with left-out data.
The optimal actions are not known, but the performance of each candidate strategy can be measured (using the predictive model), therefore the prescriptive model needs to be learned using search techniques. Standard reinforcement learning methods such as PPO and DQN are possible; the experiments in this paper use evolutionary optimization, i.e. conventional neuroevolution [35]. As in prior applications of ESP [7, 22], the network has a fixed architecture of two fully connected layers; its weights are concatenated into a vector and evolved through crossover and mutation.
5 Experiments
This section describes the main results of the project so far, including different ways to create good predictors and the evolution of effective prescriptors based on them. The behavior of the discovered prescriptors is analyzed quantitatively and through various ablation and extension studies. As a concrete implementation of the mechanisms, an interactive demo of the entire land-use recommendation system is described.
5.1 Prediction
Second, a random forest model (RF) was trained on the same data but now includes the full set of input features, in order to take advantage of RF’s ability to decide which features are most useful. The forest consisted of 100 decision trees with unrestricted maximum depth; at each split, a random subset of features was considered (where is the square root of the total number of features). To make the training time of the Global RF model feasible, it was trained with data from only 1982–2011. Preliminary experiments showed that constraining the training set in this manner did not decrease performance significantly. Both the linear regression and random forest models were trained using the scikit-learn library [26].
Third, a neural network (NeuralNet) was trained for the same task with the full set of features and data from 1851–2011 using the PyTorch library [25]. The network was fully connected, composed of an input layer, a single hidden layer of size 4096, and an output layer. The output layer took as input the concatenation of the input layer and the hidden layer in order to take advantage of the observed linear relationships more easily [4]. It was trained for three epochs with batch size 2048 using the AdamW optimization algorithm [20]. After each epoch, the learning rate was scaled by a factor of 0.1.